OpenStudy (anonymous):

If f(x) = 1 over 3x − 2 and g(x) = −2x − 3, find (f • g)(6) 15 −15 0 −17

OpenStudy (anonymous):

\[f(x)=\frac{1}{3x-2} \phantom{space} g(x)=-2x-3\] \[(f*g)(6)=f(6)*g(6)=\frac{1}{3(6)-2}*(-2(6)-3)=\frac{1}{16}*(-15)=-\frac{15}{16} \]